OpenStudy (anonymous):

I need to simplify 78125^6/7. What's the easiest way to solve this? My calculator does not have this function!

OpenStudy (akshay_budhkar):

if you want the final answer visit wolfram alfa

OpenStudy (akshay_budhkar):

http://www.wolframalpha.com/input/?i=78125^6%2F7

OpenStudy (anonymous):

find 78125^6 then take 7^root of 78125^6

OpenStudy (akshay_budhkar):

here is your answer

OpenStudy (anonymous):

Yes, I know that much, but I need to solve it myself. Can you give me a better steps to estimate the answer? I understand the process, but without a calculator with infinite digits available, I need to figure it out! I am positively stumped! They are asking for me to simplify 15625 to the 5/6 power and 4096 to the 3/4 power. I am pretty good with simple steps, but trying to calculate the large numbers runs off the calculator and I get wacky number. please try again?

OpenStudy (akshay_budhkar):

so you want to solve them without the calculator?

OpenStudy (anonymous):

I have a calculator, but it is old, only 10 digit, and doesn't have that function. Is there a way to estimate these to figure them out without a calculator?

OpenStudy (akshay_budhkar):

as far as my knowledge is concerned for the 16 years i studied theres no way but use a scientific calci.... Still if you not convinced wait for some1 to reply but there is no easier method sorry:(

OpenStudy (anonymous):

ok - can you help me with programming it in? perhaps I am not doing it correctly

OpenStudy (akshay_budhkar):

you want to program a scientific calci? or are you talking of using wolframalfa?

OpenStudy (anonymous):

I will only be allowed to use a scientific calc. no on-line access in class!

OpenStudy (anonymous):

jenn u are using which calculator give the brand and model no

OpenStudy (akshay_budhkar):

probably they won't ask you such questions if they r not allowing a scientific calci.. you can bank on that

OpenStudy (anonymous):

I don't know if it helps but \[\alpha ^{\frac{1}{n}}=\exp(\frac{1}{n}\ln( \alpha))\]

OpenStudy (anonymous):

I am using a Casio fx-115 10-digit really old!

OpenStudy (akshay_budhkar):

they wont ask jenn believe me

OpenStudy (akshay_budhkar):

or else they will allow a scientific calci

OpenStudy (anonymous):

the problem is - they did ask, and I have to solve it!

OpenStudy (anonymous):

If I use the website that was previously recommended, how do I enter the exponents on the computer keyboard? Do I use the to indicate that it's an exponent? Just want to make sure I'm doing it correctly!

OpenStudy (akshay_budhkar):

^ is used for exponent

OpenStudy (anonymous):

I will try to calculate these ridiculous problems on that site. THANK YOU so much for your help. I really appreciate it!

OpenStudy (akshay_budhkar):

you are welcome!!! don't worry we are here to help you can ask what you want

OpenStudy (akshay_budhkar):

and yes dun forget my medal lol

OpenStudy (anonymous):

What does that mean?

OpenStudy (akshay_budhkar):

you are new?

OpenStudy (anonymous):

yes!

OpenStudy (akshay_budhkar):

oh ok see a place where i answered on d right hand side can you see good answer?click on it

OpenStudy (anonymous):

do I do that for each answer? will that help?

OpenStudy (akshay_budhkar):

no for 1 answer only all the others turn on automatically

OpenStudy (akshay_budhkar):

like this

OpenStudy (anonymous):

done!

OpenStudy (akshay_budhkar):

lol yea you got it! whenever you think a person helped you just click on it..ok?

OpenStudy (akshay_budhkar):

and welcome to open study

OpenStudy (anonymous):

awesome! Thanks! I will bookmark this site and use it often!

OpenStudy (akshay_budhkar):

ya do it ppl will help you here..

OpenStudy (anonymous):

cool! Let me get back to work and I will post more questions if I get stuck! I appreciate the help!

OpenStudy (akshay_budhkar):

:)

OpenStudy (anonymous):

how do you type in a radical on the computer?